# LISA: Layerwise Importance Sampling for Memory-Efficient Large Language Model Fine-Tuning

### v2

@@ Abstract
The machine learning community has witnessed impressive advancements since the first appearance of large language models (LLMs), yet their huge memory consumption has become a major roadblock to large-scale training. Parameter Efficient Fine-Tuning techniques such as Low-Rank Adaptation (LoRA) have been proposed to alleviate this problem, but their performance still fails to match full parameter training in most large-scale fine-tuning settings. Attempting to complement this deficiency, we investigate layerwise properties of LoRA on fine-tuning tasks and observe an uncommon skewness of weight norms across different layers. Utilizing this key observation, a surprisingly simple training strategy is discovered, which outperforms both LoRA and full parameter training in a wide range of settings with memory costs as low as LoRA. We name it Layerwise Importance Sampled AdamW (LISA), a promising alternative for LoRA, which applies the idea of importance sampling to different layers in LLMs and randomly freeze most middle layers during optimization. Experimental results show that with similar or less GPU memory consumption, LISA surpasses LoRA or even full parameter tuning in downstream fine-tuning tasks, where LISA consistently outperforms LoRA by over $11\%$-$37\%$ in terms of MT-Bench scores. On large models, specifically LLaMA-2-70B, LISA achieves on-par or better performance than LoRA on MT-Bench, GSM8K, and PubMedQA, demonstrating its effectiveness across different domains.
@@
Figure 1: Training loss of LLaMA-2-7B model on Alpaca GPT-4 dataset with Full Parameter Training (FT), LoRA, and LISA.
@@ 1 Introduction
Large language models like ChatGPT excel in tasks such as writing documents, generating complex code, answering questions, and conducting human-like conversations (Ouyang et al., 2022). With LLMs being increasingly applied in diverse task domains, domain-specific fine-tuning has emerged as a key strategy to enhance their downstream capabilities (Raffel et al., 2020; Chowdhery et al., 2022; Rozière et al., 2023; OpenAI et al., 2023). Nevertheless, these methods are notably expensive, introducing major obstacles in developing large-scale models. Aiming to reduce the cost, Parameter-Efficient Fine-Tuning (PEFT) techniques, like adapter weights (Houlsby et al., 2019), prompt weights (Li and Liang, 2021), and LoRA (Hu et al., 2022) have been proposed to minimize the number of trainable parameters. Among them, LoRA is one of the most widely adopted PEFT techniques, due to its nice property of allowing the adaptor to be merged back to the base model parameters. However, LoRA’s superior performance in fine-tuning tasks has yet to reach a point that universally surpasses full parameter fine-tuning in all settings (Ding et al., 2022; Dettmers et al., 2023). In particular, it has been observed that LoRA tends to falter on large-scale datasets during continual pre-training (Lialin et al., 2023), which raises doubts about the effectiveness of LoRA under those circumstances. We attribute this to the much fewer trainable parameters of LoRA compared to the base model, which limits the representation power of LoRA training.
To overcome this shortcoming, we delve into LoRA’s training statistics in each layer, aspiring to bridge the difference between LoRA and full-parameter fine-tuning. Surprisingly, we discover that LoRA’s layerwise weight norms have an uncommonly skewed distribution, where the bottom layer and/or the top layer occupy the majority of weights during the update, while the other self-attention layers only account for a small amount, which means different layers have different importance when updating. This key observation inspires us to “sample” different layers by their importance, which exactly matches the idea of importance sampling (Kloek and Van Dijk, 1978; Zhao and Zhang, 2015).
@@
  • We discover the phenomenon of skewed weight-norm distribution across layers in both LoRA and full parameter fine-tuning, which implies the varied importance of different layers in large-scale LLM training.
  • We propose the Layerwise Importance Sampled AdamW (LISA) algorithm, a simple optimization method that is capable of scaling up to over $70$B LLMs with less or similar memory cost as LoRA.
  • We demonstrate LISA’s effectiveness in fine-tuning tasks for modern LLMs, where it outperforms LoRA by $8\%$-$36\%$ in MT-Bench and exhibits much better convergence behaviors. LISA even outperforms full parameters training under certain settings. Similar performance gain is observed across different sized models ($7$B-$70$B) and different tasks, including instruction following, medical QA, and math problems.
@@ 2.1 Large Language Models
In the realm of natural language processing (NLP), the Transformer architecture has been a revolutionary technique, initially known for its effectiveness in machine translation tasks (Vaswani et al., 2017). With the inception of models like BERT (Devlin et al., 2019) and GPT-2 (Radford et al., 2019), the approach shifted towards pre-training on extensive corpora, which led to significant performance enhancements in downstream fine-tuning tasks (Brown et al., 2020; Raffel et al., 2020; Zhang et al., 2022; Scao et al., 2022; Almazrouei et al., 2023; Touvron et al., 2023a, b; Chiang et al., 2023; Biderman et al., 2023; Jiang et al., 2024). However, the growing number of parameters in these models results in a huge GPU memory consumption, rendering the fine-tuning of large scale models ($\geq 65$B) infeasible under low resource scenarios. This has prompted a shift towards more efficient training of LLMs.
@@ 2.2 Parameter-Effieient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) methods adapt pre-trained models by fine-tuning only a subset of parameters. In general, PEFT methods can be grouped into three classes: 1) Prompt Learning methods (Hambardzumyan et al., 2021; Zhong et al., 2021; Han et al., 2021; Li and Liang, 2021; Qin and Eisner, 2021; Liu et al., 2021a; Diao et al., 2022), 2) Adapter methods (Houlsby et al., 2019; Diao et al., 2021; Hu et al., 2022; Diao et al., 2023c), and 3) Selective methods (Liu et al., 2021b, b; Li et al., 2023a). Prompt learning methods emphasize optimization of the input token or input embedding with frozen model parameters, which generally has the least training cost among all three types. Adapter methods normally introduce an auxiliary module with much fewer parameters than the original model, where updates will only be applied to the adapter module during training. Compared with them, selective methods are more closely related to LISA, which focuses on optimizing a fraction of the model’s parameters without appending extra modules. Recent advances in this domain have introduced several notable techniques through layer freezing. AutoFreeze (Liu et al., 2021b) offers an adaptive mechanism to identify layers for freezing automatically and accelerates the training process. FreezeOut (Brock et al., 2017) progressively freezes intermediate layers, resulting in significant training time reductions without notably affecting accuracy. The SmartFRZ (Li et al., 2023a) framework utilizes an attention-based predictor for layer selection, substantially cutting computation and training time while maintaining accuracy. However, none of these layer-freezing strategies has been widely adopted in the context of Large Language Models due to their inherent complexity or non-compatibility with modern memory reduction techniques (Rajbhandari et al., 2020; Rasley et al., 2020) for LLMs.
@@ 2.3 Low-Rank Adaptation (LoRA)
In contrast, the Low-Rank Adaptation (LoRA) technique is much more popular in common practices of LLM training (Hu et al., 2022). By employing low-rank matrices, LoRA reduces the number of trainable parameters, thereby lessening the computational burden and memory cost. One key strength of LoRA is its compatibility with models featuring linear layers, where the decomposed low-rank matrices can be merged back into the original model. This allows for efficient deployment without changing the model architecture. As a result, LoRA can be seamlessly combined with other techniques, such as quantization (Dettmers et al., 2023) or Mixture of Experts (Gou et al., 2023). Despite these advantages, LoRA’s performance is not universally comparable with full parameter fine-tuning. There have been tasks in Ding et al. (2022) that LoRA achieves much worse performance than full parameter training on. This phenomenon is especially evident in large-scale pre-training settings (Lialin et al., 2023), where to the best of our knowledge, only full parameter training was adopted for successful open-source LLMs (Almazrouei et al., 2023; Touvron et al., 2023a, b; Jiang et al., 2023; Zhang et al., 2024; Jiang et al., 2024).
@@ 2.4 Large-scale Optimization Algorithms
Besides approaches that change model architectures, there have also been efforts to improve the efficiency of optimization algorithms for LLMs. One branch of such is layerwise optimization, whose origin can be traced back to decades ago, where Hinton et al. (2006) pioneered an effective layer-by-layer pre-training method for Deep Belief Networks (DBN), proving the benefits of sequential layer optimization. Bengio et al. (2007) furthered this concept by showcasing the advantages of a greedy, unsupervised approach for pre-training each layer in deep networks. For large batch settings, You et al. (2017, 2019) propose LARS and LAMB, providing improvements in generalization behaviors to avoid performance degradation incurred by large batch sizes. Nevertheless, in most settings involving LLMs, Adam (Kingma and Ba, 2014; Reddi et al., 2019), and AdamW (Loshchilov and Hutter, 2017) are still the dominant optimization methods currently.
Recently, there have also been other attempts to reduce the training cost of LLMs. For example, MeZO (Malladi et al., 2023) adopted zeroth order optimization that brought significant memory savings during training. However, it also incurred a huge performance drop in multiple benchmarks, particularly in complex fine-tuning scenarios. In terms of acceleration, Sophia (Liu et al., 2023) incorporates clipped second-order information into the optimization, obtaining non-trivial speedup on LLM training. The major downsides are its intrinsic complexity of Hessian estimation and unverified empirical performance in large-size models (e.g., $\geq 65$B). In parallel to our work, Zhao et al. (2024) proposed GaLore, a memory-efficient training strategy that reduces memory cost by projecting gradients into a low-rank compact space. Yet the performance has still not surpassed full-parameter training in fine-tuning settings. To sum up, LoRA-variant methods (Hu et al., 2022; Dettmers et al., 2023; Zhao et al., 2024) with AdamW (Loshchilov and Hutter, 2017) is still the dominant paradigm for large-size LLM fine-tuning, the performance of which still demands further improvements.
@@ 3.1 Motivation
To understand how LoRA achieves effective training with only a tiny portion of parameters, we carried out empirical studies on multiple models, specially observing the weight norms across various layers. We fine-tune it on the Alpaca-GPT4 dataset (Peng et al., 2023). During the training, we meticulously recorded the mean weight norms of each layer $\ell$ at every step $t$ after updates, i.e.
@@
  • The embedding layer (wte and wpe layers for GPT2) or the language model (LM) head layer exhibited significantly larger weight norms compared to intermediary layers in LoRA, often by a factor of hundreds. This phenomenon, however, was not salient under full-parameter training settings.
Refer to captionRefer to caption
@@
This observation indicates that the update emphasis of LoRA and full parameter training differ significantly, which can be attributed to the difference in their learned knowledge. For example, in embedding layers, tokens with similar meanings, i.e. synonyms, can be projected into the same embedding space and converted to similar embeddings. LoRA may capture this similarity in language and “group” them in the low-dimension space, allowing frequent features of language meanings to be promptly identified and optimized. The price is LoRA’s limited representation power restricted by its intrinsic low-rank space. Other possible explanations can also justify this phenomenon. Despite various interpretations of this observation, one fact remains clear: LoRA values layerwise importance differently from full parameter tuning.
@@ 3.2 Layerwise Importance Sampled AdamW (LISA)
To exploit the aforementioned discovery, we aspire to simulate LoRA’s updating pattern via sampling different layers to freeze. This way, we can avoid LoRA’s inherent deficiency of limited low-rank representation ability and also emulate its fast learning process. Intuitively, given the same global learning rates across layers, layers with small weight norms in LoRA should also have small sampling probabilities to unfreeze in full-parameter settings, so the expected learning rates across iterations can stay the same. This is exactly the idea of importance sampling (Kloek and Van Dijk, 1978; Zhao and Zhang, 2015), where instead of applying layerwise different learning rates $\{\eta_{t}\}$ in full-parameter settings to emulate LoRA’s updates $\{\tilde{\eta}_{t}\}$, we apply sampling instead
@@
0:  number of layers $N_{L}$, number of iterations $T$, sampling period $K$, sampling probability $\{p_{\ell}\}_{\ell=1}^{N_{L}}$, initial learning rate $\eta_{0}$
@@
2:     for $\ell\leftarrow 1$ to $N_{L}$ do
3:        if  $\mathcal{U}(0,1)>p_{\ell}$ then
4:           Freeze layer $\ell$
5:        end if
6:     end for
7:     Run AdamW for $K$ iterations with $\{\eta_{t}\}_{t=ik}^{ik+k-1}$
8:  end for
@@ 4 Experimental Results
Table 1: The chart illustrates peak GPU memory consumption for various model architectures and configurations, highlighting differences across models. In the table, the LISA configuration is specifically labeled: “E” denotes the embedding layer, “H” represents the language model head layer, and “2L” indicates two additional standard layers. *: Model parallelism is applied for the 70B model.

### v3

@@ Abstract
The machine learning community has witnessed impressive advancements since large language models (LLMs) first appeared. Yet, their massive memory consumption has become a significant roadblock to large-scale training. For instance, a 7B model typically requires at least 60 GB of GPU memory with full parameter training, which presents challenges for researchers without access to high-resource environments. Parameter Efficient Fine-Tuning techniques such as Low-Rank Adaptation (LoRA) have been proposed to alleviate this problem. However, in most large-scale fine-tuning settings, their performance does not reach the level of full parameter training because they confine the parameter search to a low-rank subspace. Attempting to complement this deficiency, we investigate the layerwise properties of LoRA on fine-tuning tasks and observe an unexpected but consistent skewness of weight norms across different layers. Utilizing this key observation, a surprisingly simple training strategy is discovered, which outperforms both LoRA and full parameter training in a wide range of settings with memory costs as low as LoRA. We name it Layerwise Importance Sampled AdamW (LISA), a promising alternative for LoRA, which applies the idea of importance sampling to different layers in LLMs and randomly freeze most middle layers during optimization. Experimental results show that with similar or less GPU memory consumption, LISA surpasses LoRA or even full parameter tuning in downstream fine-tuning tasks, where LISA consistently outperforms LoRA by over $10\%$-$35\%$ in terms of MT-Bench score while achieving on-par or better performance in MMLU, AGIEval and WinoGrande. On large models, specifically LLaMA-2-70B, LISA surpasses LoRA on MT-Bench, GSM8K, and PubMedQA, demonstrating its effectiveness across different domains.
@@ 1 Introduction
Large language models (LLMs) like ChatGPT excel in tasks such as writing documents, generating complex code, answering questions, and conducting human-like conversations (Ouyang et al., 2022). With LLMs being increasingly applied in diverse task domains, domain-specific fine-tuning has emerged as a critical strategy to enhance their downstream capabilities (Raffel et al., 2020; Chowdhery et al., 2022; Rozière et al., 2023; OpenAI et al., 2023). Nevertheless, these methods are typically time-intensive and consume substantial computational resources, posing significant challenges to the development of large-scale models. For example, continual pre-training typically requires several weeks even with multiple 80 GB GPUs. To reduce costs, Parameter-Efficient Fine-Tuning (PEFT) techniques have been proposed to minimize the number of trainable parameters. These techniques include adapter weights (Houlsby et al., 2019), prompt weights (Li and Liang, 2021), and LoRA (Hu et al., 2022). Among these, LoRA stands out as one of the most widely adopted due to its unique ability to merge the adaptor back into the base model parameters, significantly enhancing efficiency. However, LoRA’s superior performance in fine-tuning tasks has yet to reach a point that universally surpasses full parameter fine-tuning in all settings (Ding et al., 2022; Dettmers et al., 2023). In particular, it has been observed that LoRA tends to falter on large-scale datasets during continual pre-training (Lialin et al., 2023), which raises doubts about the effectiveness of LoRA under those circumstances.
Figure 1: Training loss of LLaMA-2-7B model on Alpaca GPT-4 dataset with Full Parameter Training (FT), LoRA, GaLore, and LISA.
We attribute this to LoRA’s much fewer trainable parameters compared to the base model, which limits the representation power of LoRA training.
To overcome this shortcoming, we delve into LoRA’s training statistics in each layer, aspiring to bridge the difference between LoRA and full-parameter fine-tuning. Surprisingly, we discover that LoRA’s layerwise weight norms have an uncommonly skewed distribution, where the bottom layer and/or the top layer occupy the majority of weights during the update. In contrast, the other self-attention layers only account for a small amount, which means different layers have different importance when updating. This key observation inspires us to “sample” different layers by their importance, which matches the idea of importance sampling (Kloek and Van Dijk, 1978; Zhao and Zhang, 2015).
@@
  • We discover the phenomenon of skewed weight-norm distribution across layers in LoRA, which implies the varied importance of different layers in large-scale LLM training.
  • We propose the Layerwise Importance Sampled AdamW (LISA), a simple optimization method capable of scaling up to over $70$B LLMs with less or similar memory cost as LoRA.
  • We demonstrate LISA’s effectiveness in fine-tuning tasks for modern LLMs, where it outperforms LoRA by $10\%$-$35\%$ in MT-Bench and achieves better performance in multiple benchmarks. In addition, LISA exhibits much better convergence behaviors than LoRA. LISA even outperforms full parameters training under certain settings. Similar performance gain is observed across different sized models ($7$B-$70$B) and tasks, including instruction following, medical QA, and math problems.
@@ 2.1 Large Language Models
In the realm of natural language processing (NLP), the Transformer architecture has been a revolutionary technique, initially known for its effectiveness in machine translation tasks (Vaswani et al., 2017). With the inception of models like BERT (Devlin et al., 2019) and GPT-2 (Radford et al., 2019), the approach shifted towards pre-training on extensive corpora, which led to significant performance enhancements in downstream fine-tuning tasks (Raffel et al., 2020; Brown et al., 2020; Zhang et al., 2022; Scao et al., 2022; Almazrouei et al., 2023; Touvron et al., 2023a, b; Chiang et al., 2023; Biderman et al., 2023; Jiang et al., 2024). However, the growing number of parameters in these models results in a huge GPU memory consumption, rendering the fine-tuning of large scale models ($\geq 65$B) infeasible under low resource scenarios. This has prompted a shift towards more efficient training of LLMs.
@@ 2.2 Parameter-Effieient Fine-Tuning
Parameter-efficient fine-tuning (PEFT) methods adapt pre-trained models by fine-tuning only a subset of parameters. In general, PEFT methods can be grouped into three classes: 1) Prompt Learning methods (Li and Liang, 2021; Hambardzumyan et al., 2021; Zhong et al., 2021; Han et al., 2021; Qin and Eisner, 2021; Liu et al., 2021a; Diao et al., 2022), 2) Adapter methods (Houlsby et al., 2019; Hu et al., 2022; Diao et al., 2021, 2023c; Meng et al., 2024; SONG et al., 2024; Ji et al., 2024), and 3) Selective methods (Liu et al., 2021b, b; Li et al., 2023; Luo et al., 2024). Prompt learning methods emphasize optimizing the input token or input embedding with frozen model parameters, which generally has the least training cost among all three types. Adapter methods normally introduce an auxiliary module with much fewer parameters than the original model, and updates are only applied to the adapter module during training. Compared with them, selective methods are more closely related to LISA, which focuses on optimizing a fraction of the model’s parameters without appending extra modules. Recent advances in this domain have introduced several notable techniques through layer freezing. AutoFreeze (Liu et al., 2021b) offers an adaptive mechanism to identify layers for freezing automatically and accelerates the training process. FreezeOut (Brock et al., 2017) progressively freezes intermediate layers, significantly reducing training time without notably affecting accuracy. The SmartFRZ (Li et al., 2023) framework utilizes an attention-based predictor for layer selection, substantially cutting computation and training time while maintaining accuracy. However, none of these layer-freezing strategies has been widely adopted in the context of Large Language Models due to their inherent complexity or non-compatibility with modern memory reduction techniques (Rajbhandari et al., 2020; Rasley et al., 2020) for LLMs.
@@ 2.3 Low-Rank Adaptation (LoRA)
In contrast, the Low-Rank Adaptation (LoRA) technique is much more prevalent in common LLM training (Hu et al., 2022). LoRA reduces the number of trainable parameters by employing low-rank matrices, thereby lessening the computational burden and memory cost. One key strength of LoRA is its compatibility with models featuring linear layers, where the decomposed low-rank matrices can be merged back into the original model. This allows for efficient deployment without changing the model architecture. As a result, LoRA can be seamlessly combined with other techniques, such as quantization (Dettmers et al., 2023) or Mixture of Experts (Gou et al., 2023). Despite these advantages, LoRA’s performance is not universally comparable with full parameter fine-tuning. There have been tasks in (Ding et al., 2022) that LoRA performs much worse than full parameter training on. This phenomenon is especially evident in large-scale pre-training settings (Lialin et al., 2023), where to the best of our knowledge, only full parameter training was adopted for successful open-source LLMs (Almazrouei et al., 2023; Touvron et al., 2023a, b; Jiang et al., 2023; Zhang et al., 2024; Jiang et al., 2024).
@@ 2.4 Large-scale Optimization Algorithms
In addition to approaches that change model architectures, there have also been efforts to improve the efficiency of optimization algorithms for LLMs. One such approach is layerwise optimization, a concept with roots extending back several decades. Notably, (Hinton et al., 2006) introduced an effective layer-by-layer pre-training method for Deep Belief Networks (DBN), demonstrating the benefits of sequential layer optimization. This idea was expanded by researchers like (Bengio et al., 2007), who illustrated the advantages of a greedy, unsupervised approach to pre-training each layer of deep networks. In the context of large batch training, (You et al., 2017, 2019) developed LARS and LAMB to improve generalization and mitigate the performance declines associated with large batch sizes. Despite these innovations, Adam (Kingma and Ba, 2014; Reddi et al., 2019) and AdamW (Loshchilov and Hutter, 2017) continue to be the predominant optimization methods used in most LLM settings.
Recently, other attempts have also been made to reduce the training cost of LLMs. For example, MeZO (Malladi et al., 2023) adopted zeroth order optimization, bringing significant memory savings during training. However, it also incurred a considerable performance drop in multiple benchmarks, particularly in complex fine-tuning scenarios. Regarding acceleration, Sophia (Liu et al., 2023) incorporates clipped second-order information into the optimization, obtaining non-trivial speedup on LLM training. The significant downsides are its intrinsic complexity of Hessian estimation and unverified empirical performance in large-size models (e.g., $\geq 65$B). In parallel to our work, (Zhao et al., 2024) proposed GaLore, a memory-efficient training strategy that reduces memory cost by projecting gradients into a low-rank compact space. Yet the performance has still not surpassed full-parameter training in fine-tuning settings. To sum up, LoRA-variant methods (Hu et al., 2022; Dettmers et al., 2023; Zhao et al., 2024) with AdamW (Loshchilov and Hutter, 2017) is still the dominant paradigm for large-size LLM fine-tuning, the performance of which still demands further improvements.
@@ 3.1 Motivation
To understand how LoRA achieves effective training with only a few parameters, we conducted empirical studies on multiple models, especially observing the weight norms across various layers. We fine-tune it on the Alpaca-GPT4 dataset (Peng et al., 2023). During the training, we meticulously recorded the mean weight norms of each layer $\ell$ at every step $t$ after updates, i.e.
@@
  • The embedding layer or the language modeling (LM) head layer exhibits significantly larger weight norms than intermediary layers in LoRA, often by a factor of hundreds. This phenomenon, however, was not salient under full-parameter training settings.
@@
This observation indicates that the update emphasis of LoRA and full parameter training differ significantly, which can be attributed to the difference in their learned knowledge. For example, in embedding layers, tokens with similar meanings, i.e., synonyms, can be projected into the same embedding space and converted to similar embeddings. LoRA may capture this similarity in language and “group” them in the low-dimension space, allowing frequent features of language meanings to be promptly identified and optimized. The price is LoRA’s limited representation power restricted by its intrinsic low-rank space, as we can see from the comparison with LISA in image generation tasks (Appendix A.1), where LoRA memorizes and learns details much slower than LISA. Other possible explanations can also justify this phenomenon. Despite various interpretations of this observation, one fact remains clear: LoRA values layerwise importance differently from full parameter tuning.
@@ 3.2 Layerwise Importance Sampled AdamW (LISA)
To exploit the discovery above, we aspire to simulate LoRA’s updating pattern via sampling different layers to freeze. This way, we can avoid LoRA’s inherent deficiency of limited low-rank representation ability and emulate its fast learning process. Intuitively, given the same global learning rates across layers, layers with small weight norms in LoRA should also have small sampling probabilities to unfreeze in full-parameter settings so the expected learning rates across iterations can stay the same. This is exactly the idea of importance sampling (Kloek and Van Dijk, 1978; Zhao and Zhang, 2015), where instead of applying layerwise different learning rates $\{\eta_{t}\}$ in full-parameter settings to emulate LoRA’s updates $\{\tilde{\eta}_{t}\}$, we apply sampling and instead get the same expected parameter update
@@
0:  number of layers $N_{L}$, number of iterations $T$, sampling period $K$, number of sampled layers $\gamma$, initial learning rate $\eta_{0}$
@@
2:     Freeze all layers except the embedding and language modeling head layer
3:     Randomly sample $\gamma$ intermediate layers to unfreeze
4:     Run AdamW for $K$ iterations with $\{\eta_{t}\}_{t=ik}^{ik+k-1}$
5:  end for
@@ 4 Experimental Results
Table 1: The chart illustrates peak GPU memory consumption for various model architectures and configurations, highlighting differences across models. The LISA configuration is specifically labeled in the table: “E” denotes the embedding layer, “H” represents the language modeling head layer, and “2L” indicates two additional intermediate layers. *: Model parallelism is applied for the 70B model.